OpenStudy (anonymous):

x - 9 = 39 how do I solve this

OpenStudy (anonymous):

Is it 14.30?

OpenStudy (kropot72):

If you add 9 to the left hand side of the equation it will cancel the -9, leaving just x on the left hand side. If you add 9 to the right hand side of the equation you get what? 39 + 9 = ?

OpenStudy (anonymous):

I have a 5th grade level in math trying to help my sophomore daughter. ..I am unsure at how to do this..is it 48?

OpenStudy (kropot72):

Correct. So we solved it as follows: x - 9 = 39 Adding 9 to both sides gives: x - 9 + 9 = 39 + 9 Simplifying we get: x = 48

OpenStudy (anonymous):

Ty so very much...I really appreciate it! You helped me understand:) ty for your patience

OpenStudy (kropot72):

Remember, whatever operation you do to one side of an equation you must also do to the other side. This keeps the equation balanced. I added 9 to both sides to eliminate the -9 and get x by itself on the left hand side.